Duncan takes a break from studying and goes to the gym to swim laps. If swimming burns 7.25×105 cal per hour, how many kilojoules does swimming burn in that same amount of time?
In other words, convert 7.25×105 cal to kilojoules.
1
Expert's answer
2019-09-12T05:30:36-0400

As 1 cal = 4.184 × 10-3 kJ:

7.25 × 105 cal = 7.25 ×105 cal × 4.184 10-3 kJ = 30.334 × 102 kJ = 3033.4 kJ


Answer: 3033.4 kilojoules